OpenStudy (akgorgeous):

The first four terms of a sequence are shown below: 7, 4, 1, −2 Which of the following functions best defines this sequence? f(1) = 7, f(n + 1) = f(n) + 3; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) − 4; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) + 4; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) − 3; for n ≥ 1

OpenStudy (tkhunny):

Normally, this is a substitution. Pick n = 1 and see what happens, f(1 + 1) = f(2) = f(1) + 3 = 7 + 3 = 10 f(1 + 1) = f(2) = f(1) − 4 = 7 - 4 = 3 f(1 + 1) = f(2) = f(1) + 4 = 7 + 4 = 11 f(1 + 1) = f(2) = f(1) − 3 = 7 - 3 = 4 Which, if any, actually reproduced the second value?

OpenStudy (cp123):

D